OpenStudy (anonymous):

-x = -36 x = __________ 4x - 6 = 34 x = __________ -5x - 7 = 108 x = __________

OpenStudy (anonymous):

-x = -36 x = 36 4x - 6 = 34 x = 10 -5x - 7 = 108 x = -23

OpenStudy (anonymous):

1st one- divide both sides by -1 2nd one- add 6 to both sides then divide by 4 3rd one- add 7 to both side then divide by -5